Calculate the angular velocity and linear velocity of a tip of minute hand of length 10 cm. 

Grade:12th pass

1 Answers

SAHIL
7 years ago
Given : T = 60 min. = 60 × 60 s = 3600 s
L = 10 cm = 0.1 m
To Find : ω = ?
v = ?
Formula : ω = π T 2 v = rω
Solution : ω = π T 2 = 2 3.142 × 3600
∴ ω = 1.744 × 10–3 rad/s
 
v = rω = 0.1 × 1.745 × 10–3 ∴ v = 1.745 × 10–4 m/s